OpenStudy (anonymous):

How can the equation x + 2 = 5 be solved using algebra tiles? A. add two 1-tiles to each side B. remove two 1-tiles from the left side C. remove two 1-tiles from each side D. remove five 1-tiles from each side

OpenStudy (anonymous):

I answered this here for you sirstaven. :) http://openstudy.com/study#/updates/53cd26fde4b0f2f7e5e650ac

OpenStudy (anonymous):

Should I ask my next 2 questions here?

OpenStudy (anonymous):

If the solution of an equation is x = –2, what could the original equation be? A. x + 6 = 8 B. x – 6 = 8 C. x + 8 = 6 D. x – 8 = 6